Multiple choice questions
Question 5
Because of the nature of the power flow bus data, nodal equations do not directly fit
the linear-equation format, and power flow equations are actually nonlinear. However,
the Gauss-Siedel method can be used for the power flow solution.
(1) True.
(2) Only true for systems with less than 14 buses.
(3) False.
(4) Only true for systems with fully transposed transmission lines.
1
Expert's answer
2021-08-09T05:11:01-0400

The given statement is True (Option 1)

Consider the following for a bus k

Pk = real component of power

Qk = reactive component of power

Vk = Voltage magnitude

Since power-flow bus data consists of Pk and Qk for load buses or Pk and Vk for voltage-controlled buses, nodal equations do not directly fit the linear equation format and due to this reason, the current source vector I is unknown and the equations are actually nonlinear. The Gauss-Siedel method can be applied for the power-flow equation in the following way: 

"v_k (i+1)= \frac{1}{Y_{kk}}[\frac{P_k-jQ_k}{V^*_k(i)}-\sum_{n=1}^{k-1}Y_{kn}V_n(i+1)-\sum_{n=k+1}^{N}Y_{kn}V_n(i)]"

Where, Ykn is the element in the kth row and nth column of matrix Y. Hence, the given statement is 

True.